# Eye-gaze Guided Multi-modal Alignment Framework for Radiology

### v2

@@ Appendix A Supplementary Materials
The supplementary document is organized as follows:
  • Additional Experiments: In Sec. B, we provide additional experiments of zero-shot classification task after continue pre-training using the backbones of other SOTA models in our EGMA framework.
  • Experimental Details: In Sec. C, we provide more experimental settings, including training parameters, detailed parameters of image and text encoders.
  • Dataset Descriptions: In Sec. D, we introduce the detailed information of the datasets used in this work.
  • Potential Negative Impact: In Sec. E, we discuss the potential negative impacts of this work.
Appendix B Additional Experiments
To further validate the generality of our framework, we utilize the encoders and pre-training weights provided by CLIP [1], GLoRIA [10], and MGCA [11] in our EGMA framework. Subsequently, we continue training on the MIMIC-EYE [14] dataset and present the zero-shot classification results on the CheXpert 5x200 [6], RSNA [36], and SIIM-ACR [37] datasets in Tab. IV.
TABLE IV: Comparison results of zero-shot classification after continue pre-training using the backbones of other SOTA models in our EGMA framework. Bold denotes the best result and the values in (parentheses) represents the improvement over the baseline metrics in Table 1 of main manuscript.
Method CheXpert 5x200 [6] RSNA [36] SIIM-ACR [37]
EGMA(Ours) 60.50 58.44 57.81 36.70 63.62 61.46
We present accuracy Accuracy and F1 score metrics on three datasets. The values in parentheses indicate the improvement over the baseline metrics (as shown in Table 1 of main manuscript). It can be observed that, all models show improvement after training with our EGMA framework, except for the decrease in metrics for the trained MGCA model on the RSNA dataset. For MGCA, when tested with its provided pre-trained weights, it performs the best on the RSNA dataset (as shown in Table 1 of main manuscript), but after training with the EGMA framework, its performance improves on CheXpert 5x200 and SIIM and decreases on RSNA dataset. This may reflect that the features extracted by MGCA on RSNA dataset are not truly disease-related features but rather shortcut features, indicating that the high baseline metrics were based on easily distinguishable shortcut features. Furthermore, after training with EGMA, the performance of other models significantly improves on all three datasets.
Appendix C Experimental Details
C-A Training Settings
In our study, all chest X-ray images and eye-gaze heatmaps are padded to square then scaled to $224\times 224$. In the training process, we utilize the following image augmentations to the chest X-ray images: scale to images to $224\times 224$; color jittering with brightness and contrast ratios from [0.8, 1.2]; randomly change the contrast($probability=0.5$). And our batch size is 100.
C-B Image/Text Encoder
@@
Similar to [9], we use SwinTransformer [38] as the image encoder, BioClinicalBERT [39] as the text encoder. Specifically, we use a 4 stages SwinTransformer, including 2, 2, 6, and 2 SwinTransformer blocks. Other parameters are: patch size 4; window size 7. And we use a 6 layers BioClinicalBERT with 12 attention heads. In our EGMA framework, we add a linear projection layer after both the image encoder and text encoder to map the embeddings’ dimension to 512, and we use a learnable temperature $\tau$ in contrastive loss calculation initialized on 0.07.
Appendix D Dataset Descriptions
D-A MIMIC-EYE
@@
The MIMIC-EYE [14] dataset includes a comprehensive range of patient information, including medical images and reports, clinical data, patient’s hospital journey, and eye-tracking data and audio of radiologists during diagnosis. The dataset comprises a total of 3689 images from the MIMIC-IV v1.0 dataset [31], each accompanied by transcripts text from audio and eye-tracking data of radiologists. In this work, we use this dataset as our training set. As described in Figure 3 of our main manuscript, we first segment the transcribed text at the sentence level. After determining the start and end time points for each sentence, we obtain eye-gaze information corresponding to the respective time intervals. Subsequently, we generate sentence-level heatmaps, which are then fed into our EGMA framework for training.
D-B CheXpert
@@
severity increased
mild improved small
minimal apperance of stable
@@ D-D SIIM-ACR
Appendix E Potential Negative Impact
Our EGMA framework explores the feasibility of incorporating eye-gaze data from radiologists to assist in multi-modal feature alignment during model training, laying the foundation for the application of eye-gaze data in the medical domain. Although the eye-gaze data we used is publicly available and we have permission to use it, some studies [47, 48] have indicated that private information such as gender, age, and mental state of observers can be extracted from eye-gaze data. Therefore, privacy concerns have always been a focal point in eye-gaze data collection. Our suggestion is not to public the raw eye-gaze data but to release heatmaps generated from gaze points, significantly reducing the privacy risks associated with the data while still enabling its application in subsequent training strategy.

### v3

@@ Appendix A Supplementary Materials
The supplementary document is organized as follows. In Sec. B, we provide more experimental settings, including training parameters, detailed parameters of image and text encoders. In Sec. C, we introduce the detailed information of the datasets used in this work. In Sec. D, we provide more details of multi-modal data processing of MIMIC-EYE [33] dataset. In Sec. E, we provide additional visualization results of feature representation. In Sec. F, we provide additional experiments of zero-shot classification task after continue pre-training using the backbones of other SOTA models in our EGMA framework.
Appendix B Experimental Details
B-A Image/Text Encoder
@@
In this study, we use SwinTransformer [47] as the image encoder, BioClinicalBERT [48] as the text encoder. Specifically, we use a 4 stages SwinTransformer, including 2, 2, 6, and 2 SwinTransformer blocks. Other parameters are: patch size 4; window size 7. And we use a 6 layers BioClinicalBERT with 12 attention heads. In our EGMA framework, we add a linear projection layer after both the image encoder and text encoder to map the embeddings’ dimension to 512, and we use a learnable temperature $\tau$ in contrastive loss calculation initialized on 0.07.
B-B Training Settings
Pre-training Settings In the pre-training process, we utilize the following image augmentations to the chest X-ray images: scale to images to $224\times 224$; color jittering with brightness and contrast ratios from [0.8, 1.2]; randomly change the contrast($probability=0.5$). And we train our model with 50 epochs with an initial learning rate $1\times 10^{-6}$ and weight decay $1\times 10^{-4}$ and 10 epochs of warm-up.
Fine-tuning Settings In the supervised classification experiments, we adopt the linear classification settings [10], where the pre-trained image encoder is frozen, and only a randomly initialized linear classification head is trained. We choose the same image augmentations to the above pre-training settings. And we fine-tune our model with 30 epochs with an initial learning rate $5\times 10^{-7}$ and weight decay $1\times 10^{-4}$ and 6 epochs of warm-up. And all our training tasks are completed on four RTX 3090 GPUs.
Appendix C Dataset Descriptions
C-A MIMIC-EYE
@@
The MIMIC-EYE [33] dataset includes a comprehensive range of patient information, including medical images and reports, clinical data, patient’s hospital journey, and eye-tracking data and audio of radiologists during diagnosis. The dataset comprises a total of 3689 images from the MIMIC-IV v1.0 dataset [36], each accompanied by transcripts text from audio and eye-tracking data of radiologists. In this work, we use this dataset as our training set.
C-B CheXpert
@@
severity mild increased small
minimal improved stable
apperance of
@@ C-D SIIM-ACR
Appendix D Details of Multi-modal Data Processing
As illustrated in Fig. 5, the presentation of multi-modal data in the MIMIC-EYE [33] dataset includes radiologists’ audio, text transcript, eye-gaze data, and image. Since each modality is synchronized, the audio data is aligned with the eye-gaze data in time. By segmenting the audio based on the time before and after the pronunciation of each word, we can align the transcripts with the audio, thereby aligning word-level text with eye-gaze data. Subsequently, we generate attention heatmap based on eye-gaze data and images to represent the image regions the radiologist focuses on. Through the aforementioned data processing steps, we achieve precise alignment between word-level text and image regions. It is noteworthy that due to the rapid speech rate of radiologists, there may be no available eye-gaze data within the time interval corresponding to a single word. In Fig. 5, the word “with” in the transcript has no corresponding gaze data. Another common and unavoidable issue is the loss of eye-gaze data caused by blinking and intense head movement of radiologist, as seen in the last two words of the transcript. Due to these technical challenges, achieving perfect pairing between words and image regions is difficult. However, as shown in the right side of Fig. 5, adjusting the text to the sentence level largely mitigates the issue of missing word-level heatmap (Heatmap with red edge), and the semantic information of the entire sentence also encompasses the information of each word. Therefore, in this work, we process text features at the sentence level.
During the pre-training of EGMA, the size of the input heatmap is determined by the number of patches in the image. For example, after the image is processed by the image encoder, the size of image embedding is $196\times 768$, where 196 represents the number of image patches. Therefore, we resize the heatmap directly to $14\times 14$ to match the image embedding and further process it into the Gaze-guided Similarity and Gaze-guided Label mentioned in the main manuscript.
Appendix E Additional Visualization Results
TABLE VII: Comparison results of zero-shot classification after continue pre-training using the backbones of other SOTA models in our EGMA framework. Red and blue denote the best and second-best results. The values in (parentheses) represents the improvement over the baseline metrics in Table 1 of main manuscript.
Method CheXpert 5x200 [39] RSNA [40] SIIM-ACR [41]
EGMA(Ours) 61.30 60.38 76.97 43.49 63.62 61.46
Appendix F Additional Analysis Results
To further validate the generality of our framework, we utilize the encoders and pre-training weights provided by CLIP [1], GLoRIA [10], and MGCA [11] in our EGMA framework. Subsequently, we continue training on the MIMIC-EYE [33] dataset and present the zero-shot classification results on the CheXpert 5x200 [39], RSNA [40], and SIIM-ACR [41] datasets in Tab. VII.
We present accuracy Accuracy and F1 score metrics on three datasets. The values in parentheses indicate the improvement over the baseline metrics (as shown in Tab. II). It can be observed that, all models show improvement after training with our EGMA framework, except for the decrease in metrics for the trained MGCA model on the RSNA dataset. For MGCA, when tested with its provided pre-trained weights, it performs the best F1-score on the RSNA dataset (as shown in Tab. II), but after training with the EGMA framework, its performance improves on CheXpert 5x200 and SIIM but decreases on RSNA dataset. This may reflect that the features extracted by MGCA on RSNA dataset are not truly disease-related features but rather shortcut features, indicating that the high baseline metrics were based on easily distinguishable shortcut features. Furthermore, after training with EGMA, the performance of other models significantly improves on all three datasets.